OpenStudy (anonymous):

4y + xy = 2 find the second derivative.

OpenStudy (anonymous):

at (-3,2)

OpenStudy (shamil98):

Compute the first derivative first, what do you get?

OpenStudy (anonymous):

How do you find the derivative of this. It doesn't look like the derivatives I usually do.

OpenStudy (shamil98):

Implicit differentiation

OpenStudy (shamil98):

differentiate with respect to x, though, do you know how to do implicit differentiation?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You can either proceed by using implicit differentiation or notting that \[\large 4y+xy = 2\implies y(4+x)=2\implies y=\frac{2}{4+x}\] and then differentiate like you normally would (using chain rule, of course).

OpenStudy (anonymous):

I know how to do normal differentiation. Is it different?

OpenStudy (shamil98):

\[\huge \frac{ d }{ dx } 4y + \frac{ d }{ dx } xy = \frac{ d }{ dx}2\] Use the product rule for xy btw. it's a bit different

OpenStudy (shamil98):

\[\huge 4y' + y + xy' = 0\] Get the y' on one side and the other terms on the other. \[\huge y'(4+x) = -y\] divide \[\huge y' = \frac{ -y }{ 4+x }\]

OpenStudy (anonymous):

Okay and then do I use the quotient rule to find the second derivative?

OpenStudy (shamil98):

yeah.

OpenStudy (anonymous):

okay one minute

OpenStudy (anonymous):

the derivative of -y would just be 0 right?

OpenStudy (anonymous):